Statistics show that a certain soccer player has a 63% chance of missing the goal each time he shoots. If this player shoots twice, the exact probability that he scores a goal both times is
If he misses 63%, then he makes it 37% each time he shoots.
To calculate the probability if he makes it both times, you multiply it together:
37% * 37% = 13.69%
Answer:
The probability of scoring two goals in both times is 0.137 or 13.7%
Step-by-step explanation:
Statistics show that a certain soccer player has a 63% chance of missing the goal each time he shoots.
So \(P(A)=0.63\)
Hence, the probability of getting success in his shoots will be,
\(P(A^c)=1-P(A)=1-0.63=0.37\)
The probability of scoring two goals in both times is,
\(=P(A^c)\cdot P(A^c)\)
\(=0.37\times 0.37\)
\(=0.137\)
\(=13.7\%\)